How many ounces of pure gold must be added to five ounces of something that is 1% gold in order that the mixture is 6% gold?
I tried working it out but I am not sure if I did it right.
x + 5(.01)= .06(x+5)
x + .05 = .06x + .30
x + .05 - .05 - .06x = .06x - .06x + .30 - .05
.94x = .25
.94x/.94 = .25/.94
x= .27 ounces

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

How many ounces of pure gold must be added to five ounces of something that is 1% gold in order that the mixture is 6% gold?
x + 5(.01)= .06(x+5)
x + .05 = .06x + .30
x + .05 - .05 - .06x = .06x - .06x + .30 - .05
.94x = .25
.94x/.94 = .25/.94
x= .27 ounces
================
Correct Nice job.
Cheers,
Stan H.

Answer by edjones(8007)

Let x=oz of pure gold
amount of alloy=5*.01= .05
amount of pure gold=1.00*x= x
amount of pure gold + alloy= .06(5+x)=.3+.06x
.05+x=.3+.06x
.94x=.25
x=.266 oz of pure gold
Check:
.266+5=5.266 total oz
.05+.266=.316 oz of pure gold
.316/5.266=.06 or 6%
You were right.
It was more difficult than usual to chek this one.
Keep up the good work!
Ed